# Plasma Mixing in Collisionless Kelvin–Helmholtz Dynamics

### v1

@@ Appendix A Semi-Implicit PIC Methods
The results shown in this Letter correspond to the nonlinear stage of the Kelvin–Helmholtz instability. We start the simulation with the IMM until the KHI develops, and then switch to ECSIM at the early stages of the nonlinear phase, so that all results presented here are obtained using ECSIM.
@@ Appendix B Effect of the In-plane Magnetic Field on Plasma Mixing
Panels (a) and (c) of Figure 4 show the spatial distribution of the percentage of mixed plasma, $\tilde{n}/n_{0}$, for ions and electrons, respectively, in the presence of a finite $B_{x}$. In this case, as KH vortices develop, plasma mixing remains weak and is largely confined to the narrow interface between the two plasma regions. Although vortical structures form and advect plasma across the shear layer, the in-plane magnetic field limits cross-field transport, resulting in well-defined vortices and localized mixing (up to $30\%$) primarily concentrated near the frontier, in line with the discussion presented in the main text. In contrast, panels (b) and (d) show the corresponding ion and electron distributions, always for the lower shear in the absence of an in-plane magnetic field. In this case, the nonlinear evolution of the KHI is accompanied by significantly enhanced plasma mixing (up to $50\%$), with particles transported far from the initial interface and mixing regions extending well into both plasma domains. The absence of magnetic tension allows vortical motions to more efficiently redistribute plasma across the shear layer, leading to a broader and more spatially extended mixing region.

### v2

@@ Appendix A Semi-Implicit PIC Methods
The results shown in this work correspond to the nonlinear stage of the Kelvin–Helmholtz instability. We start the simulation with the IMM until the KHI develops, and then switch to ECSIM at the early stages of the nonlinear phase, so that all results presented here are obtained using the ECSIM method.
@@ Appendix B Effect of the In-plane Magnetic Field on Plasma Mixing
Panels (a) and (c) of Figure 4 show the spatial distribution of the percentage of mixed plasma, $\tilde{n}/n_{0}$, for ions and electrons, respectively, in the presence of a finite $B_{x}$. In this case, as KH vortices develop, plasma mixing remains weak and is largely confined to the narrow interface between the two plasma regions. Although vortical structures form and advect plasma across the shear layer, the in-plane magnetic field limits cross-field transport, resulting in well-defined vortices and localized mixing (up to $30\%$) primarily concentrated near the frontier, in line with the discussion presented in the main text. In contrast, panels (b) and (d) show the corresponding ion and electron distributions, always for the lower shear, this time in the absence of an in-plane magnetic field. In this case, the nonlinear evolution of the KHI is accompanied by significantly enhanced plasma mixing (up to $50\%$), with particles transported far from the initial interface and mixing regions extending well into both plasma domains. The absence of magnetic tension allows vortical motions to more efficiently redistribute plasma across the shear layer, leading to a broader and more spatially extended mixing region.